Contents
- 1 What is the probability of rolling a 12 on a die?
- 2 How many outcomes are possible from tossing a coin and rolling a 12 sided die?
- 3 What is the probability of rolling a 9 on a 12 sided die?
- 4 Is rolling a 12 sided die a uniform distribution?
- 5 What is the probability of rolling a 4?
- 6 How many times must you roll a die until each side has?
- 7 How many numbers can you roll on two dice in craps?
- 8 Which is the only number that can be rolled more than once?
What is the probability of rolling a 12 on a die?
Probabilities for the two dice
| Total | Number of combinations | Probability |
|---|---|---|
| 10 | 3 | 8.33% |
| 11 | 2 | 5.56% |
| 12 | 1 | 2.78% |
| Total | 36 | 100% |
How many outcomes are possible from tossing a coin and rolling a 12 sided die?
SOLUTION: List each outcome from tossing a coin. Then pair each coin outcome with an outcome from rolling a 12-sided solid. There are 24 possible outcomes.
How many possibilities are there for rolling a die?
36 possibilities
Note that there are 36 possibilities for (a,b). This total number of possibilities can be obtained from the multiplication principle: there are 6 possibilities for a, and for each outcome for a, there are 6 possibilities for b. So, the total number of joint outcomes (a,b) is 6 times 6 which is 36.
What is the probability of rolling a 9 on a 12 sided die?
The odds of rolling a 9 is 1 in 12. The odds of rolling a 9 or a 12 is 2 in 12 or 1 in 6.
Is rolling a 12 sided die a uniform distribution?
In the scenario of rolling a fair die or tossing a fair coin, each simple event is uniformly distributed, and its random variable always approaches a uniform distribution.
What is the probability of flipping a tails and rolling a 4?
1/16
1) Consider the experiment of flipping of 4 coins. If we assume that each individual coin is equally likely to come up heads or tails, then each of the above 16 outcomes to 4 flips is equally likely. Each occurs a fraction one out of 16 times, or each has a probability of 1/16.
What is the probability of rolling a 4?
Two (6-sided) dice roll probability table
| Roll a… | Probability |
|---|---|
| 4 | 6/36 (16.667%) |
| 5 | 10/36 (27.778%) |
| 6 | 15/36 (41.667%) |
| 7 | 21/36 (58.333%) |
How many times must you roll a die until each side has?
After that value appears, the probability of rolling a new face is 4/6, and therefore you expect the third face after 6/4 rolls. Continuing this process leads to the conclusion that the expected number of rolls before all six faces appear is 6/6 + 6/5 + 6/4 + 6/3 + 6/2 + 6/1 = 14.7 rolls.
How to calculate the probability of rolling a die?
The probability of rolling the same value on each die – while the chance of getting a particular value on a single die is p, we only need to multiply this probability by itself as many times as the number of dice. In other words, the probability P equals p to the power n, or P = pⁿ = (1/s)ⁿ.
How many numbers can you roll on two dice in craps?
It follows that the number of possible combinations is equal to 6 x 6 = 36 as each of the two dice used in craps has six sides. Each throw of the two dice will result in one of these eleven numbers coming out. Some numbers, however, are rolled more frequently as the number of combinations that add up to them is greater.
Which is the only number that can be rolled more than once?
Some numbers, however, are rolled more frequently as the number of combinations that add up to them is greater. Thus, there is only one possible combination for numbers 2 (1 plus 1) and 12 (6 plus 6) as opposed to number 7, where 3 plus 4, 1 plus 6, 5 plus 2 and their respective permutations can all add up to its total.